# KernelSHAP-IQ: Weighted Least Square Optimization for Shapley Interactions

### v1

@@ 4 Experiments
Baselines. We compare KernelSHAP-IQ and inconsistent KernelSHAP-IQ with all available baseline algorithms; permutation sampling (Castro et al., 2009; Tsai et al., 2023), SHAP-IQ (Fumagalli et al., 2023), and SVARM-IQ (Kolpaczki et al., 2024), as shown in Table 1.
@@
Benchmark Datasets and Models. Based on recent work by Fumagalli et al. (2023); Kolpaczki et al. (2024); Tsai et al. (2023), we create different benchmark scenarios222github.com/FFmgll/KernelSHAP_IQ. Table 2 summarizes the scenarios and corresponding removal approaches after Covert et al. (2021). For a detailed description regarding the experimental setup and model training we refer to Appendices C and D including a runtime analysis. First, we create synthetic sum of unanimity models (SOUMs), also known as sum of unanimity games and induced subgraph game (Deng & Papadimitriou, 1994), where GT SII can be computed, cf. Section B.8. Second, we compute feature (i.e. token) interaction values for a sentiment analysis language model (LM). The LM is a fine-tuned version of DistilBert (Sanh et al., 2019) on movie review excerpts from the IMDB dataset (Maas et al., 2011). Third and forth, we explain local instances of the bike rental (BR) (Fanaee-T & Gama, 2014) and california housing (CH) (Kelley Pace & Barry, 1997) regression datasets. On BR we train an XGBoost regressor and on CH a small neural network. The target variables are logarithmized. Fifth and sixth, we explain a ResNet18 convolutional neural network (CNN) (He et al., 2016) and a vision transformer (ViT) image classifiers, which were pre-trained on ImageNet (Deng et al., 2009). The ViT considers patches of $32\times 32$ pixels. In the CNN, individual pixels are grouped together into super-pixels. Seventh, we compute feature interactions for instances of the adult census (AC) (Kohavi, 1996) classification dataset and a random forest (RF) classifier.
@@
Approximation Quality of SII and $k$-SII. Figure 4 depicts the approximation quality of KernelSHAP-IQ and inconsistent KernelSHAP-IQ compared to the baseline estimators on a selection of benchmark tasks. Further results are provided in Appendix D. KernelSHAP-IQ and inconsistent KernelSHAP-IQ clearly outperform all three sampling-based baselines on both regression benchmarks (BR and CH). Inconsistent KernelSHAP-IQ achieves high estimation qualities in low-budget scenarios. This trend also materializes for second order SII values, in particular in high-player tasks (e.g. SOUM in Figure 4). Yet, unlike KernelSHAP-IQ, the inconsistent version does not converge to the GT SII values. In contrast to Pelegrina et al. (2023), we observe that the inconsistent version does converge to the GT SV independent of the approximation order $k\geq 1$. Notably, in benchmark tasks like the LM, KernelSHAP-IQ and SVARM-IQ rapidly outperform all baselines as well as inconsistent KernelSHAP-IQ. Both SVARM-IQ and KernelSHAP-IQ achieve SOTA results for the LM task and second order SII values. According to 3.9, KernelSHAP-IQ yields high-quality estimations for SII of higher orders ($k>2$) and outperforms SVARM-IQ for order $3$ on the LM.
@@
Example Use Case: Feature Attribution. As illustrated in Figures 1 and 2, 2-SII values can be used to enhance current feature attribution techniques. Similar to Fumagalli et al. (2023); Sundararajan et al. (2020); Tsai et al. (2023), we show that feature interactions are relevant for understanding intricate LMs. In Figure 1, 2-SII scores reveal that the predicted positive sentiment largely stems from the interaction of the two words “never” and “forget”, while “forget” individually points towards a negative sentiment. In the CH example illustrated in Figure 2, the longit. and latit. features are both contributing positively to the prediction considering 1-SII. The 2-SII explanation, however, reveals that the positive contribution of latit. vanishes (very low latit. score) and can be mostly attributed to the latit. x longit. interaction. Hence, the exact location of the property is meaningful.
@@ 6 Conclusion and Future Work
In future work, we suspect that a rigorous proof of our conjectures would reveal novel techniques and further insights. Another interesting link are orthogonal projections, where pairwise SII can be viewed as a decomposition of the Shapley residual (Kumar et al., 2021). Our novel representation of SII opens up additional possibilities, including the investigation of amortized Shapley interactions through methods akin to FastSHAP (Jethani et al., 2022). Apart from local interpretability with feature interactions, KernelSHAP-IQ is applicable in any game-theoretic setting, which includes global interpretability or data valuation.
@@ Acknowledgment
Fabian Fumagalli and Maximilan Muschalik gratefully acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation): TRR 318/1 2021 – 438445824. Patrick Kolpaczki was supported supported by the research training group Dataninja (Trustworthy AI for Seamless Problem Solving: Next Generation Intelligence Joins Robust Data Analysis) funded by the German federal state of North Rhine-Westphalia.
@@
This paper presents work whose goal is to advance the field of Machine Learning, specifically the field of Explainable Artificial Intelligence. There are many potential societal consequences of our work. Our work can positively impact Machine Learning adoption and potentially reveal biases or unwanted behaviour in Machine Learning system. However, it may also enable explainability-based white-washing in organizations, firms, or policing.
@@ Appendix A Proofs
By Proposition 3.2, the $n$-order approximation is given as
$\hat{\nu}_{n}(T)=\sum_{S\subseteq N}^{1\leq|S|\leq n}\phi^{\text{SII}}(S)\lambda(|S|,|S\cap T|)=\sum_{S\subseteq N}^{1\leq|S|\leq n}\phi^{\text{SII}}(S)\left(\sum_{r=1}^{|T\cap S|}\binom{|T\cap S|}{r}B_{|S|-|T\cap S|}\right),$
which yields $\nu(T)=\hat{\nu}_{n}(T)$ by a well-known result from cooperative game theory (Denneberg & Grabisch, 1999), cf. Table 4 in (Grabisch et al., 2000). ∎
@@
Note that for $T\in\{\emptyset,N\}$, we have $(\mathbf{y}_{1})_{T}=\nu(T)-\hat{\nu}_{1}(T)=\nu(T)-\sum_{i\in T}\phi^{\text{SV}}(i)=0$ and thus we do not have to consider these cases. The proof is again structured as follows:
@@
=(11)${=}-\frac{1}{2}\mu_{\infty}((n-1)\frac{\tau^{\uparrow}}{\tau^{\downarrow}}+n-3)(\tilde{\mu}_{2,0}-\tilde{\mu}_{2,1})$ )
@@ B.1 KernelSHAP-IQ Default Parameters
    We further use similar sampling weights for the baseline methods, if applicable, i.e. SHAP-IQ (Fumagalli et al., 2023) and SVARM-IQ (Kolpaczki et al., 2024).
  • Border-Trick: We use the border-trick (Fumagalli et al., 2023; Lundberg & Lee, 2017) to split the sampling procedure in a deterministic and a sampling part, cf. Section B.3. We also apply this method to the baseline algorithms, were it is applicable, i.e. SHAP-IQ (Fumagalli et al., 2023) and SVARM-IQ (Kolpaczki et al., 2024).
@@ B.2 Inconsistent KernelSHAP-IQ
5:  end for
6:  $(\hat{\mathbf{W}}^{*}_{\leq k})_{TT}\leftarrow\mu_{1}(t)\cdot w_{T}$triangleright weight adjustment
@@
8:  $(\mathbf{Z}_{k})_{S\tilde{S}}\leftarrow\textsc{{kSIIWeight}}(S,\tilde{S})$ for $1\leq|S|,|\tilde{S}|\leq k$
9:  $\hat{\Phi}_{k}\leftarrow\mathbf{Z}_{k}\cdot\hat{\phi}_{\leq k}$triangleright aggregate SII to $k$-SII
@@ B.7 Baseline Methods: SHAP-IQ, Permutation Sampling and SVARM-IQ
The pseudocode and implementations can be found in the corresponding paper. Permutation sampling (Tsai et al., 2023; Fumagalli et al., 2023) is described for SII as an extension to permutation sampling of the SV (Castro et al., 2009). SHAP-IQ (Fumagalli et al., 2023) is implemented using similar sampling weights as KernelSHAP-IQ and extends Unbiased KernelSHAP (Covert & Lee, 2021) to interactions. SVARM-IQ (Kolpaczki et al., 2024) extends SVARM (Kolpaczki et al., 2023) to interactions.
@@ B.8 Analytic Solution for SII of SOUMs
$\phi^{\text{SII}}(S)$ $=\sum_{T\subseteq N\setminus S}\frac{(n-t-s)!t!}{(n-s+1)!}\Delta_{S}(T)=\mathbf{1}_{S\subseteq R}\sum_{T\subseteq N\setminus S}\frac{(n-t-s)!t!}{(n-s+1)!}\mathbf{1}_{R\setminus S\subseteq T}=\mathbf{1}_{S\subseteq R}\sum_{t=q}^{n-s}\frac{(n-t-s)!t!}{(n-s+1)!}\binom{n-s-q}{t-q}$
$=\mathbf{1}_{S\subseteq R}\sum_{t=0}^{n-s-q}\frac{(t+q)!}{(n-s+1)!}\frac{(n-s-q)!}{t!}=\mathbf{1}_{S\subseteq R}\frac{(n-s-q)!}{(n-s+1)!}\sum_{t=0}^{n-s-q}\frac{(t+q)!}{q!}=\mathbf{1}_{S\subseteq R}\frac{(n-s-q)!}{(n-s+1)!}q!\sum_{t=0}^{n-s-q}\binom{t+q}{q}$
$=\mathbf{1}_{S\subseteq R}\frac{(n-s-q)!}{(n-s+1)!}q!\sum_{t=0}^{n-s-q}\binom{t+q}{q}=\mathbf{1}_{S\subseteq R}\frac{(n-s-q)!}{(n-s+1)!}q!\binom{n-s+1}{n-s-q}=\mathbf{1}_{S\subseteq R}\frac{1}{q+1}=\frac{\mathbf{1}_{S\subseteq R}}{r-s+1},$
where we used the hockey-stick identity in the last row to compute the sum and $q=r-s$, if $S\subseteq R$. ∎
@@ Appendix C Experimental Setup and Reproducibility
This section contains additional information about the setup and reproducibility of our empirical evaluation. Section C.1 contains additional information about the models, datasets and explanation tasks used for evaluation as summarized in Table 2. Section C.2 contains additional information about the environmental impact and computational effort of the empirical evaluation. All experiments can be reproduced via the technical supplement available at github.com/FFmgll/KernelSHAP_IQwhich will be made fully publicly available.
@@ C.1 Model, Datasets and Task Descriptions
This section contains detailed information about the models and datasets used for benchmarking the approximation quality of KernelSHAP-IQ and the available baseline algorithms. To increase the comparability, we conduct our empirical evaluation of the approximation quality on settings and tasks presented in the literature (Kolpaczki et al., 2024; Fumagalli et al., 2023; Muschalik et al., 2024; Sundararajan et al., 2020; Tsai et al., 2023). For further questions regarding the technical details, we further refer to the technical supplement at github.com/FFmgll/KernelSHAP_IQwhich will be made fully publicly available.
@@ C.1.1 Sum Of Unanimity Models (SOUM)
The sum of unanimity models (SOUMs) are used in different empirical evaluations of Shapley-based interaction estimators (Fumagalli et al., 2023; Kolpaczki et al., 2024; Tsai et al., 2023). This synthetic model class can be used to create explanation tasks with differing complexity while allowing to compute GT interaction scores analytically even for higher player counts. For a given player set $N:=\{1,\dots,n\}$ with $n$ many players, we sample $M=50$ interactions $R_{1},\ldots,R_{M}\subseteq N$ from $\mathcal{P}(N)$. Next, we draw for each interaction subset $R_{m}$ a coefficient $a_{m}\in[0,1]$ uniformly at random. The value function is then defined as $\nu(T)=\sum_{m=1}^{M}a_{m}\cdot\nu_{R_{m}}(T)$ for all coalitions $T\subseteq N$ with the unanimity game $\nu_{R_{m}}(T):=\mathbf{1}_{T\supseteq R_{m}}$ Similar to Tsai et al. (2023), we limit the highest interaction size $|R_{m}|\leq 4$. Further, we randomly set two features to be non-informative (dummy player) which are never part of any interaction subset.

### v2

@@ 4 Experiments
Baselines. We compare KernelSHAP-IQ and inconsistent KernelSHAP-IQ with all available baseline algorithms; permutation sampling (Castro et al., 2009; Tsai et al., 2023), SHAP-IQ (Fumagalli et al., 2023), and SVARM-IQ (Kolpaczki et al., 2024b), as shown in Table 1.
@@
Benchmark Datasets and Models. Based on recent work by Fumagalli et al. (2023); Kolpaczki et al. (2024b); Tsai et al. (2023), we create different benchmark scenarios [^2] https://github.com/FFmgll/KernelSHAP_IQ_Supplementary_Material. Table 2 summarizes the scenarios and corresponding removal approaches after Covert et al. (2021). For a detailed description regarding the experimental setup and model training we refer to Appendices C and D including a runtime analysis. First, we create synthetic sum of unanimity models (SOUMs), also known as sum of unanimity games and induced subgraph game (Deng & Papadimitriou, 1994), where GT SII can be computed, cf. Section B.8. Second, we compute feature (i.e. token) interaction values for a sentiment analysis language model (LM). The LM is a fine-tuned version of DistilBert (Sanh et al., 2019) on movie review excerpts from the IMDB dataset (Maas et al., 2011). Third and forth, we explain local instances of the bike rental (BR) (Fanaee-T & Gama, 2014) and california housing (CH) (Kelley Pace & Barry, 1997) regression datasets. On BR we train an XGBoost regressor and on CH a small neural network. The target variables are logarithmized. Fifth and sixth, we explain a ResNet18 convolutional neural network (CNN) (He et al., 2016) and a vision transformer (ViT) image classifiers, which were pre-trained on ImageNet (Deng et al., 2009). The ViT considers patches of $32\times 32$ pixels. In the CNN, individual pixels are grouped together into super-pixels. Seventh, we compute feature interactions for instances of the adult census (AC) (Kohavi, 1996) classification dataset and a random forest (RF) classifier.
@@
Approximation Quality of SII and $k$-SII. Figure 4 depicts the approximation quality of KernelSHAP-IQ and inconsistent KernelSHAP-IQ compared to the baseline estimators on a selection of benchmark tasks. Further results are provided in Appendix D. KernelSHAP-IQ and inconsistent KernelSHAP-IQ clearly outperform all three sampling-based baselines on both regression benchmarks (BR and CH). Inconsistent KernelSHAP-IQ achieves high estimation qualities in low-budget scenarios. This trend also materializes for second order SII values, in particular in high-player tasks (e.g. SOUM in Figure 4). Yet, unlike KernelSHAP-IQ, the inconsistent version does not converge to the GT SII values. Interestingly, we observe that the inconsistent version does converge to the GT SV independent of the approximation order $k\geq 1$. Notably, in benchmark tasks like the LM, KernelSHAP-IQ and SVARM-IQ rapidly outperform all baselines as well as inconsistent KernelSHAP-IQ. Both SVARM-IQ and KernelSHAP-IQ achieve SOTA results for the LM task and second order SII values. By relying on 3.9, KernelSHAP-IQ yields high-quality estimations for SII of higher orders ($k>2$) and outperforms SVARM-IQ for order $3$ on the LM.
@@
Example Use Case: Feature Attribution. As illustrated in Figures 1 and 2, 2-SII values can be used to enhance current feature attribution techniques. Similar to Fumagalli et al. (2023); Sundararajan et al. (2020); Tsai et al. (2023), we show that feature interactions are relevant for understanding intricate LMs. In Figure 1, 2-SII scores reveal that the predicted positive sentiment largely stems from the interaction of the two words “never” and “forget”, while “forget” individually points towards a negative sentiment. In the CH example illustrated in Figure 2, the longit. and latit. features are both contributing positively to the prediction considering 1-SII (SV). The 2-SII explanation, however, reveals that the positive contribution of latit. vanishes (very low latit. score) and can be mostly attributed to the latit. x longit. interaction. Hence, the exact location of the property is meaningful.
@@ 6 Conclusion and Future Work
In future work, we suspect that a rigorous proof of our conjectures would reveal novel techniques and further insights. Another interesting link are orthogonal projections, where we suspect a link of pairwise SII to the Shapley residual (Kumar et al., 2021). Our novel representation of SII opens up additional possibilities, including the investigation of amortized Shapley interactions through methods akin to FastSHAP (Jethani et al., 2022). Apart from local interpretability with feature interactions, KernelSHAP-IQ is applicable in any game-theoretic setting, which includes global interpretability or data valuation.
@@ Acknowledgment
We gratefully thank the anonymous reviewer for their valuable feedback for improving this work. Fabian Fumagalli and Maximilian Muschalik gratefully acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation): TRR 318/1 2021 – 438445824. Patrick Kolpaczki was supported supported by the research training group Dataninja (Trustworthy AI for Seamless Problem Solving: Next Generation Intelligence Joins Robust Data Analysis) funded by the German federal state of North Rhine-Westphalia.
@@
This paper presents work whose goal is to advance the field of Machine Learning, specifically the field of Explainable Artificial Intelligence. There are many potential societal consequences of our work. Our work can positively impact Machine Learning adoption and potentially reveal biases or unwanted behavior in Machine Learning systems. However, it may also enable explainability-based white-washing in organizations, firms, or policing.
@@ Appendix A Proofs
This proof follows immediately from the fact that $\hat{\nu}_{n}(T)$ is constructed from $k$-SII values $\Phi_{k}$, which are the Möbius transform for $k=n$, cf. Theorem 4 by Bordt & von Luxburg (2023). A defining property of the Möbius transform (Grabisch, 2016) is
$\nu(T)=\sum_{S\subseteq T}\Phi_{n}(S),$
and thus, by Definition 3.1, $\nu(T)=\hat{\nu}_{n}(T)$. ∎
@@
Note that for $T\in\{\emptyset,N\}$, we have $(\mathbf{y}_{2})_{T}=\nu(T)-\hat{\nu}_{1}(T)=\nu(T)-\sum_{i\in T}\phi^{\text{SV}}(i)=0$ and thus we do not have to consider these cases. The proof is again structured as follows:
@@
=(${=}-\frac{1}{2}\mu_{\infty}((n-1)\frac{\tau^{\uparrow}}{\tau^{\downarrow}}+n-3)(\tilde{\mu}_{2,0}-\tilde{\mu}_{2,1})$ )
@@ B.1 KernelSHAP-IQ Default Parameters
    We further use similar sampling weights for the baseline methods, if applicable, i.e. SHAP-IQ (Fumagalli et al., 2023) and SVARM-IQ (Kolpaczki et al., 2024b).
  • Border-Trick: We use the border-trick (Fumagalli et al., 2023; Lundberg & Lee, 2017) to split the sampling procedure in a deterministic and a sampling part, cf. Section B.3. We also apply this method to the baseline algorithms, were it is applicable, i.e. SHAP-IQ (Fumagalli et al., 2023) and SVARM-IQ (Kolpaczki et al., 2024b).
@@ B.2 Inconsistent KernelSHAP-IQ
5:     $(\hat{\mathbf{W}}^{*}_{\leq k})_{TT}\leftarrow\mu_{1}(t)\cdot w_{T}$triangleright weight adjustment
6:  end for
@@
8:  $\hat{\phi}_{1},\dots,\hat{\phi}_{k}\leftarrow\textsc{{Unstack}}(\hat{\phi}_{\leq k})$
9:  $\hat{\Phi}_{k}\leftarrow\textsc{{AggregateSII}}$$(\hat{\phi}_{1},\dots,\hat{\phi}_{k})$triangleright compute $k$-SII
@@ B.7 Baseline Methods: SHAP-IQ, Permutation Sampling and SVARM-IQ
The pseudocode and implementations can be found in the corresponding paper. Permutation sampling (Tsai et al., 2023; Fumagalli et al., 2023) is described for SII as an extension to permutation sampling of the SV (Castro et al., 2009). SHAP-IQ (Fumagalli et al., 2023) is implemented using similar sampling weights as KernelSHAP-IQ and extends Unbiased KernelSHAP (Covert & Lee, 2021) to interactions. SVARM-IQ (Kolpaczki et al., 2024b) extends SVARM (Kolpaczki et al., 2024a) to interactions.
@@ B.8 Analytic Solution for SII of SOUMs
$\phi^{\text{SII}}(S)$ $=\sum_{T\subseteq N\setminus S}\frac{(n-t-s)!t!}{(n-s+1)!}\Delta_{S}(T)$ $=\mathbf{1}_{S\subseteq R}\sum_{T\subseteq N\setminus S}\frac{(n-t-s)!t!}{(n-s+1)!}\mathbf{1}_{R\setminus S\subseteq T}$
$=\mathbf{1}_{S\subseteq R}\sum_{t=q}^{n-s}\frac{(n-t-s)!t!}{(n-s+1)!}\binom{n-s-q}{t-q}$ $=\mathbf{1}_{S\subseteq R}\sum_{t=0}^{n-s-q}\frac{(t+q)!}{(n-s+1)!}\frac{(n-s-q)!}{t!}$
$=\mathbf{1}_{S\subseteq R}\frac{(n-s-q)!}{(n-s+1)!}\sum_{t=0}^{n-s-q}\frac{(t+q)!}{q!}$ $=\mathbf{1}_{S\subseteq R}\frac{(n-s-q)!}{(n-s+1)!}q!\sum_{t=0}^{n-s-q}\binom{t+q}{q}$
$=\mathbf{1}_{S\subseteq R}\frac{(n-s-q)!}{(n-s+1)!}q!\sum_{t=0}^{n-s-q}\binom{t+q}{q}$ $=\mathbf{1}_{S\subseteq R}\frac{(n-s-q)!}{(n-s+1)!}q!\binom{n-s+1}{n-s-q}$
$=\mathbf{1}_{S\subseteq R}\frac{1}{q+1}$ $=\frac{\mathbf{1}_{S\subseteq R}}{r-s+1},$
where we used the hockey-stick identity in the second last row to compute the sum and $q=r-s$, if $S\subseteq R$. ∎
B.9 Intuition about the KernelSHAP-IQ Weights
The intuition behind $\mathbf{W}_{k}$ is based on common factors in two representations of SII: One being the solution of the WLS problem from Equation 3
$\phi_{k}^{\text{SII}}=(\mathbf{X}^{T}_{k}\mathbf{W}_{k}\mathbf{X}_{k})^{-1}\mathbf{X}_{k}^{T}\mathbf{W}_{k}\cdot\mathbf{y}_{k},$
and the other being the representation of 3.9 and by Fumagalli et al. (2023)
$\phi_{k}^{\text{SII}}=\mathbf{Q}_{k}\cdot\mathbf{y}_{k}.$
In detail, if we consider the final weight of a single subset $T\subseteq N$ with $k\leq|T|\leq n-k$ and all interactions $S\subseteq N$ of order $|S|=k$, then the common factors for all $S$ are according to the $S$-row and $T$-column of $\mathbf{Q}_{k}$, i.e. $(-1)^{k-|T\cap S|}m_{k}(t-|T\cap S|)$, determined by the common factors in $m_{k}(t-0),\dots,m_{k}(t-k)$. By definition of $m_{k}(t)=\frac{(n-k-t)!t!}{(n-k+1)!}$, we have $m_{k}(t-\ell)=\frac{(n-k-t+\ell)!(t-\ell)!}{(n-k+1)!}$ and thus the common factors in $\mathbf{Q}_{k}$ for all $S$-rows given a $T$-column are $\frac{(n-k-t)!(t-k)!}{(n-k+1)!}\propto\binom{n-2k}{t-k}^{-1}$. On the other hand, by Equation 3, the matrix $(\mathbf{X}^{T}_{k}\mathbf{W}_{k}\mathbf{X}_{k})^{-1}$ is independent of the subset $T$, and hence $\mu_{k}(t)$, as it contains one row and one column for each interaction $S$ of order $k$. This implies
$((\mathbf{X}^{T}_{k}\mathbf{W}_{k}\mathbf{X}_{k})^{-1}\mathbf{X}^{T}_{k}\mathbf{W}_{k})_{ST}\propto\mu_{k}(t)$
for all interactions $S$ of order $k$. Combining both proportionality results, we conclude that $\mu_{k}(t)\propto\binom{n-2k}{t-k}^{-1}$ is a suitable candidate. Note that this reasoning works for subsets with $k\leq|T|\leq n-k$.
@@ Appendix C Experimental Setup and Reproducibility
This section contains additional information about the setup and reproducibility of our empirical evaluation. Section C.1 contains additional information about the models, datasets and explanation tasks used for evaluation as summarized in Table 2. Section C.2 contains additional information about the environmental impact and computational effort of the empirical evaluation. All experiments can be reproduced via the technical supplement available at https://github.com/FFmgll/KernelSHAP_IQ_Supplementary_Materialwhich will be made fully publicly available.
@@ C.1 Model, Datasets and Task Descriptions
This section contains detailed information about the models and datasets used for benchmarking the approximation quality of KernelSHAP-IQ and the available baseline algorithms. To increase the comparability, we conduct our empirical evaluation of the approximation quality on settings and tasks presented in the literature (Kolpaczki et al., 2024b; Fumagalli et al., 2023; Muschalik et al., 2024; Sundararajan et al., 2020; Tsai et al., 2023). For further questions regarding the technical details, we further refer to the technical supplement at https://github.com/FFmgll/KernelSHAP_IQ_Supplementary_Materialwhich will be made fully publicly available.
@@ C.1.1 Sum Of Unanimity Models (SOUM)
The sum of unanimity models (SOUMs) are used in different empirical evaluations of Shapley-based interaction estimators (Fumagalli et al., 2023; Kolpaczki et al., 2024b; Tsai et al., 2023). This synthetic model class can be used to create explanation tasks with differing complexity while allowing to compute GT interaction scores analytically even for higher player counts. For a given player set $N:=\{1,\dots,n\}$ with $n$ many players, we sample $M=50$ interactions $R_{1},\ldots,R_{M}\subseteq N$ from $\mathcal{P}(N)$. Next, we draw for each interaction subset $R_{m}$ a coefficient $a_{m}\in[0,1]$ uniformly at random. The value function is then defined as $\nu(T)=\sum_{m=1}^{M}a_{m}\cdot\nu_{R_{m}}(T)$ for all coalitions $T\subseteq N$ with the unanimity game $\nu_{R_{m}}(T):=\mathbf{1}_{T\supseteq R_{m}}$ Similar to Tsai et al. (2023), we limit the highest interaction size $|R_{m}|\leq 4$. Further, we randomly set two features to be non-informative (dummy player) which are never part of any interaction subset.